OpenStudy (anonymous):

If triangle LMN is similar to triangle XYZ, which of the following is not necessarily true? angle L is congruent to angle X angle M is congruent to angle Y length of side M N over length of side N L equals length of side Y Z over length of side Z X length of side L M over length of side M N equals length of side X Y over length of side X Z.

OpenStudy (anonymous):

what do you know @JAXXDP and i will help

OpenStudy (anonymous):

Given: \[\triangle LMN \sim \triangle XYZ\] \[\huge \rightarrow \angle L \cong \angle X\] \[\huge \rightarrow \angle M \cong \angle Y\] \[\huge \rightarrow \angle N \cong \angle Z\] and \[\frac{ M N}{YZ}=\frac{NL}{Z X} \](coresponding sides of similar triangles) i.e. \[\frac{ M N}{N L}=\frac{Y Z}{Z X}\] (by alternendo) All the above are true but since MN and XZ are not corresponding sides hence the last statement "length of side L M over length of side M N equals length of side X Y over length of side X Z." is not true.